A 1500 kg van with a velocity of 22 m/s 

is going east when it hits a solid concrete 

wall in 11 s. What is the net force applied on 

the car that stopped it


1
Expert's answer
2021-06-03T09:17:46-0400

Find the deceleration that caused the van to fully stop from 22 m/s in 11 s:


"a=\frac{\Delta v}{t}=\frac{v-v_0}{t}=\frac{v}{t}."

Force, according to Newton's second law:


"F=ma=\frac{mv}{t}=3000\text{ N}."